OpenStudy (anonymous):

A balloon rising 15sec/ft is 200 feet high when a car traveling 66ft/sec passes directly beneath it. How fast is the distance between the car and the balloon increasing 1 second later?

OpenStudy (lgbasallote):

is this calculus?

OpenStudy (anonymous):

yep

OpenStudy (lgbasallote):

aw sorry i hate that part of calc =))))

OpenStudy (mathmate):

Is it 15sec/ft or 15ft/s?

OpenStudy (anonymous):

ft/sec

OpenStudy (mathmate):

Use the chain rule: Distance, D=sqrt((h^2+x^2)) differentiate with respect to t: dD/dt=dD/dh*dh/dt + dD/dx dx/dt Both dx/dt and dh/dt are known. dD/dh=2h/sqrt(h^2+x^2) and dD/dx=2x/sqrt(h^2+x^2)

OpenStudy (anonymous):

can you show me how to use the rule for this problem?

OpenStudy (mathmate):

First draw a tree diagram of what function is dependent on what. We know that the distance D is dependent on h and x. In turn, x is dependent on t, and h as well. So if we need to differentiate D with respect to t, we musnt't forget the relationship of h and x with t, so dD/dt=dD/dx * dx/dt +... according to the chain rule Since D is also dependent on h, we do the same thing, but add the effects of each, so dD/dt=dD/dh*dh/dt + dD/dx dx/dt dh/dt=15 fps dx/dt=66mph=? fps and dD/dx and dD/dh have been given above, so you can take it from here?

OpenStudy (phi):

 solve for dr/dt. You know x and dx/dt, y and dy/dt. you can find r (at the given time, 1 sec after the car passes beneath the balloon)